# Knowledge-aware Text-Image Retrieval for Remote Sensing Images

### v1

@@ V-B Results on Remote Sensing Benchmarks
Figure 7: Examples of KTIR’s and the baseline model (BLIP)’s top 5 retrieval results on the UCM-Caption dataset for both text-image retrieval and image-text retrieval. Texts or images in green boxes are predictions that correspond to the ground truth. If the Texts or images are in yellow boxes, these results are not annotated as the ground truth but are regarded as reasonable predictions according to human evaluation. Texts or images in red are semantically incorrect retrieval results.
@@
UCM-Caption (Table III)
RSICD (Table IV)
RSITMD (Table V)
@@ V-C Qualitative Analysis
Some visual examples of retrieval are shown in Fig. 7: KTIR not only has higher retrieval performance but also orders the candidate captions or images better in the ranking. In the text-image examples, the images retrieved are in a better order while using external knowledge. Similarly, in the image-text example, integrating external information redirects the model’s attention towards a more prominent part of the image.
We also perform an open-set text-image retrieval experiment, where the captions are not seen by the model during training. Results are shown in Fig. 8. By integrating external knowledge, the model can generalize well to unseen captions. When encountering unseen concepts (e.g., sunshade and walk), KTIR links them with known objects based on commonsense knowledge while the baseline fails to retrieve reasonable results.
@@ V-D Ablations and Parameter Analysis
We design ablation studies to provide further insights into the proposed KTIR method, including the analysis of the model architecture and the loss functions. The results are shown in Table VI. Starting from the image encoder ($\operatorname{ViT}$) and text-only mode of text encoder ($\operatorname{BERT}_{text}$) with knowledge-aware contrastive loss ($\mathcal{L}_{con}$) (row 1), we add the multimodal mode of the text encoder ($\operatorname{BERT}_{multi}$) and the knowledge-aware matching loss ($\mathcal{L}_{mat}$) (row 2). Then we add the cross-attention layer to fuse the captions and knowledge (row 3). The impact of the cross-attention layer can be seen in row 2 and row 3. By using a cross-attention mechanism to fuse the knowledge embedding and text embedding, the model slightly improves its performance (0.55% on mR). The model gains 1.05% by using two losses with respect to the contrastive objective only indicating knowledge emphasizes the effectiveness of the text-image matching objective proposed in BLIP. This may be due to the lack of scene and description diversity in the specific dataset. In this case, adding additional pairing information helps model training.
TABLE VI: Ablation studies of model architecture and loss functions on the UCM-Caption dataset. CrossAtt denotes the cross-attention layer to fuse the knowledge embedding and the caption embedding.
@@
TABLE VII: Parameter analysis of the number of triplets on UCM-Caption dataset. “$m$” demotes the number of triplets per caption.
@@
The number of triplets ($m$) that are used during retrieval is an important parameter of KTIR. We set $m$ to 1, 3, and 5 to study the impact of this parameter on the UCM-Caption results. We randomly select $m$ triplets from the triplet lists to generate the knowledge sentences if the total number of triplets is more than $m$, otherwise, we keep the original number of the triplets. The results show that adding more additional information to the text will further help the model to align the two modalities. When $m=5$, mR improves 2.22% compared to $m=1$ on the image-text retrieval ($i2t$) and 0.98% overall. The opposite behaviour is observed in the text-image retrieval ($t2i$). As the number of triplets rises, the effectiveness of text-image retrieval diminishes. This suggests that as text representations become more varied, the task of retrieving the same image from various texts becomes more challenging, whereas discerning text based on images becomes simpler. Note that we exclude $m>5$, as in such cases, the knowledge sentences may consistently exceed the length of the captions.
V-E Backbone Analysis

### v2

@@ V-B Results on Remote Sensing Benchmarks
UCM-Caption (Table IV)
RSICD (Table V)
Figure 7: Examples of KTIR’s and the baseline model (BLIP)’s top 5 retrieval results on the UCM-Caption dataset for both text-image retrieval and image-text retrieval. Texts or images in green boxes are predictions that correspond to the ground truth. If the Texts or images are in yellow boxes, these results are not annotated as the ground truth but are regarded as reasonable predictions according to human evaluation. Texts or images in red are semantically incorrect retrieval results. Note that retrieved images or texts matching all the semantic concepts are considered reasonable results, while results missing one or more semantic concepts are considered incorrect results.
@@
RSITMD (Table VI)
@@ V-C Qualitative Analysis
Some visual examples of retrieval are shown in Fig. 7: KTIR not only has higher retrieval performance but also orders the candidate captions or images better in the ranking. In the text-image examples, the images retrieved are more reasonable. Similarly, in the image-text example, integrating external information redirects the model’s attention towards a more prominent part of the image.
We also perform an open-set text-image retrieval experiment, where some concepts in the caption are not seen by the model during training. Results are shown in Fig. 8. By integrating external knowledge, the model can generalize well to unseen captions. When encountering unseen concepts (e.g., sunshade and walk), KTIR links them with known objects based on commonsense knowledge while the baseline fails to retrieve reasonable results.
@@ V-D Ablations and Parameter Analysis
We design ablation studies to provide further insights into the proposed KTIR method, including the analysis of the model architecture and the loss functions. The results are shown in Table VII. Starting from the image encoder ($\operatorname{ViT}$) and text-only mode of text encoder ($\operatorname{BERT}_{text}$) with knowledge-aware contrastive loss ($\mathcal{L}_{con}$) (row 1), we add the multimodal mode of the text encoder ($\operatorname{BERT}_{multi}$) and the knowledge-aware matching loss ($\mathcal{L}_{mat}$) (row 2). Then we add the cross-attention layer to fuse the captions and knowledge (row 3). The impact of the cross-attention layer can be seen in row 2 and row 3. By using a cross-attention mechanism to fuse the knowledge embedding and text embedding, the model slightly improves its performance (0.55% on mR). The model gains 1.05% by using two losses with respect to the contrastive objective only indicating knowledge emphasizes the effectiveness of the text-image matching objective proposed in BLIP. This may be due to the lack of scene and description diversity in the specific dataset. In this case, adding additional pairing information helps model training.
TABLE VII: Ablation studies of model architecture and loss functions on the UCM-Caption dataset. CrossAtt denotes the cross-attention layer to fuse the knowledge embedding and the caption embedding. $\mathrm{mR}_{t2i}$ and $\mathrm{mR}_{i2t}$ denote the average of R@$k$ in text-image and image-text retrieval, respectively.
@@
TABLE VIII: Parameter analysis of different loss weight settings on UCM-Caption dataset. “$w_{1}$” and “$w_{2}$” denote the weights of two loss terms in Eq. 10.
$w_{1}$ ($\mathcal{L}_{con}$) $w_{2}$ ($\mathcal{L}_{mat}$) mRt2i mRi2t mR
$0.5$ $1$ 59.97 56.67 58.32
$1$ $0.5$ 59.24 58.25 58.75
$1$ $1$ 59.90 57.62 58.76
TABLE IX: Parameter analysis of the number of triplets on UCM-Caption dataset. “$m$” demotes the number of triplets per caption.
@@
$m=7$ 59.46 57.46 58.46
$m=10$ 59.58 55.87 57.73
To study the impact of the loss weights (Eq. 10), we tested the model with different values of $(w_{1}$, $w_{2})$ : $(0.5,1)$, $(1,0.5)$ and $(1,1)$. The results in Table VIII suggest that the best performance is achieved when the two losses have the same weights ($w_{1}=w_{2}=1$). When $w_{1}>w_{2}$, the importance of the contrastive loss component increases, leading to a 0.63% improvement in image-text retrieval performance. This indicates the impact of the contrastive objective in retrieving the right text based on the image content. When $w_{2}>w_{1}$, the performance of text-image retrieval improves instead (mRt2i increases 0.07%). This suggests that the matching objective plays a significant role in retrieving images based on text descriptions. These findings also align with the trends observed in the loss ablation results (see rows $1$ and $2$ in Table VII).
The number of triplets ($m$) that are used during retrieval is an important parameter of KTIR: it controls the proportion of information extracted from external knowledge sentences and fed into the final text. We varied $m$ to 1, 3, 5, 7 and 10 to study the impact of this parameter on the UCM-Caption dataset in Table IX. We randomly select $m$ triplets from the triplet lists to generate the knowledge sentences if the total number of triplets is more than $m$, otherwise, we keep the original number of the triplets. The results show that the balance between original information and external knowledge is achieved when $m=5$. The mR improves 2.22% compared to $m=1$ on the image-text retrieval ($i2t$) and 0.98% overall. The opposite behaviour is observed in the text-image retrieval ($t2i$). As the number of triplets rises from 1 to 5, the effectiveness of text-image retrieval decreases. This suggests that as text representations become more varied, the task of retrieving the same image from various texts becomes more challenging, whereas discerning text based on images becomes simpler. When $m>5$, the performance of image-text retrieval drops more dramatically than in the case of text-image retrieval. This is likely due to the fact that, as the number of knowledge sentences increases, the information from the caption contributes less to the final text representation and distinguishing text descriptions becomes more challenging.
V-E Triplet selection methods
TABLE X: Experimental results on the UCM-Caption dataset with different triplet selection methods.
$m$ mRt2i mRi2t mR
Relevance with the captions 59.05 57.30 58.18
Diversity among triplets 59.56 56.51 58.04
Random selection 59.90 57.62 58.76
Selecting knowledge triplets at random could be sub-optimal. To understand the impact of different triplet selection methods, we tested KTIR with three approaches:
  • Relevance to the caption. We first measure the semantic similarity between the triplets and their corresponding caption using a pre-trained sentence BERT model [50]. To maximize the use of relevant triplets to expand the caption’s content, we then rank the triplets in reverse order based on the similarity scores and select the top $m$ ones, i.e., those with the largest semantic distance from the caption.
  • Diversity among triplets. We randomly select the first triplet, and then measure the semantic relevance score between the remaining triplets and the selected one. To maximize diversity among the selected triplets, we then rank the remaining ones in reverse order based on the semantic relevance scores and choose the top $m$ ones.
  • Random selection. We shuffle the retrieved triplet list and randomly select $m$ triplets from all the possible ones.
We use the above triplet selection methods to select triplets and train the model separately on the UCM-Caption dataset. The experimental results are reported in Table X. Although relevance is the basis of the knowledge-aware method to retrieve and extract information from knowledge sources, we observe that an active selection of the knowledge triplet does not seem to improve the results numerically: random selection yields the best overall performance (gains of on average 0.65% in terms of mR). On one hand, the success of random selection may be attributed to the fact that the exploration of the knowledge graph is limited since knowledge retrieval is restricted to the one-step neighborhood on the graph. Therefore, even with random selection, the knowledge triplets maintain high semantic relevance to the captions. On the other hand, random selection might introduce more triplet combinations and higher error tolerance to the retrieval process which could also benefit the overall effectiveness. The results also encourage future research on strategies to explore knowledge sources that balance semantic relevance and information diversity.
V-F Backbone Analysis